I need some help
Given f(x)=3x^3-4x-6, find vertex,domain, range and axis of symmetry

Thanks for your time

Answer by stanbon(75887)

Given f(x)=3x^2-4x-6, find vertex,domain, range and axis of symmetry
I assume you meant 3x^2 and not 3x^3
-----------------------------
Rewrite as:
y=3(x^2-(4/3)x-2)
x^2-(4/3)x-2=(1/3)y
x^2-(4/3)x+? = (1/3)(y+6)+?
Complete the square:
x^2-(4/3)x+(2/3)^2 = (1/3)(y+6)+(4/9)
(x-(2/3))^2 = (1/3)(y+6+(4/3))
(x-(2/3))^2 = (1/3)(y+(22/3))
Vertex is at x=2/3; y=-22/3
Domain is all real numbers.
Range is all y>=-22/3
Axis of symmetry is x=2/3
Cheers,
Stan H.